U=F(R,S) and r=x+y , s=x-y show that (∂u/∂x)+(∂u/∂y)=2(∂u/∂r)

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
Write RHS du/dx+du/dy= du/dr*(dr/dx+dr/dy)
Now dr/dx= 1 and dr/dy= 1
So RHS= du/dr*(1+1)= 2du/dr= LHS
hence proved